OpenStudy (anonymous):

Find the Period of \[2 cos^3 x + 5cos4x\]

OpenStudy (anonymous):

\(\cos^3x\)'s period is \(2\pi\) and \(\cos 4x\)'s period is \(\pi/2\). The period for the given equations should be \(2\pi\).

OpenStudy (anonymous):

equation*

OpenStudy (anonymous):

lol....u seems to find the answer easily

OpenStudy (anonymous):

any trick behind that

OpenStudy (anonymous):

Practice. :-)

OpenStudy (anonymous):

in my...book the soln is given half a page

OpenStudy (anonymous):

@Ishaan94 hw cos^3 x period is 2 pie

OpenStudy (anonymous):

Try plotting \(\cos^3x\). You will see it yourself or you can prove it as well.

OpenStudy (anonymous):

Is this +2 's topic

OpenStudy (anonymous):

idk. but you it's pretty basic you can learn it even class 10th.

OpenStudy (anonymous):

you know* ...

OpenStudy (anonymous):

lol...not in 10 th....

OpenStudy (anonymous):

yes. you can. :-)

OpenStudy (anonymous):

thxxxx